# MAESTRO: Multi-Agent Environment Shaping through Task and Reward Optimization

### v1

@@ 8 Conclusion
We introduced MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), a contextual LLM-aided MARL framework that repositions the LLM from a real-time decision-maker to an offline training architect. MAESTRO integrates an LLM-driven adaptive curriculum with LLM-generated reward functions, while leaving the underlying MADDPG policy unchanged during deployment. This separation preserves low-latency inference while leveraging the LLM’s semantic reasoning capability to shape the training process.
[…]
Overall, our findings highlight that LLMs are particularly effective as architects of the training process. When used to generate structured curricula and constrained reward functions, they complement efficient domain-specific MARL learners and enable robust policy learning in complex, non-stationary control environments.
@@ A.1 Per-Seed Performance Breakdown
Table 3 provides complete per-seed performance data over episodes 180–199. A8 exhibits the widest performance range (11.36 points), with seed 400 achieving the best result (166.08) but seed 500 falling below baseline (154.72). In contrast, A7’s two completed seeds show remarkable consistency (range: 1.69 points), both exceeding baseline.
Table 3: Per-seed traffic control performance (episode return over episodes 180–199). Training rewards extracted from simulation logs. A7 has 2 completed seeds; A2 and A8 have 4 seeds each.
@@
A7’s worst-case performance (157.92) exceeds the baseline mean (156.93) by 0.99 points, while A8’s worst-case (154.72) falls 2.21 points below baseline. This represents a critical difference: A7 guarantees improvement regardless of random seed, whereas A8 carries 25% risk of degradation.
@@
  • A2: $\pm$[1.91, 0.18, 2.49, 4.24] points (max deviation: 4.24)
  • A7: $\pm$[0.85, 0.84] points (max deviation: 0.85)
@@ A.2 Curriculum Adaptation Dynamics
Figure 7: Curriculum difficulty trajectories over 200 episodes for representative seeds. Both A2 and A7 (sharing llm_adaptive mode) exhibit bidirectional adaptation: difficulty increases when performance improves but decreases after performance drops. Shaded regions indicate episodes where difficulty was reduced to prevent overchallenge.
@@ A.4 Computational Cost Breakdown
For a full 4-seed ablation study: $\sim$\$1.88 (A7 with 2 seeds) + \$1.28 (A2/A8 with 4 seeds each) = \$3.16 total.
@@ A.5 Traffic Metrics: Multi-Seed Analysis
Table 6: Traffic efficiency metrics extracted from simulation logs (late-training episodes, seed 200). Future work will aggregate across all seeds for stronger statistical support.
@@
Current analysis uses seed 200 only due to time constraints. Future work will extract and aggregate traffic metrics across all seeds (seeds 200, 300, 400, 500 for A2/A8; seeds 200, 300 for A7) to provide mean $\pm$ SD for each metric, strengthening statistical rigor.
@@ A.7 Additional Risk Metrics
A2 (Baseline) 152.69 4.24 75% (3/4)
A7 (LLM Rewards) 1.53 2.16 157.92 0.85 100% (2/2)
A8 (Alt Curriculum) 0.70 1.12 154.72 5.51 75% (3/4)
@@ A.8 Configuration Details
  • Reward: LLM-generated with adaptive weighting, scaled 170.0, $w(d)\in[0.1,0.5]$
  • Seeds: 200, 300 (completed); 400, 500 (incomplete, stopped at episode 16)

### v2

@@ 8 Conclusion
We introduced MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), a contextual LLM-aided MARL framework that uses an LLM as an offline training architect rather than an online controller. By combining an adaptive semantic curriculum with template-based LLM reward shaping and prior policy regularization on top of MADDPG, MAESTRO improves traffic efficiency on a realistic 16-intersection network while preserving low-latency deployment.
A controlled ablation across three LLM-assisted configurations yields two core insights: (i) the LLM’s curriculum generation is the dominant lever for performance, and (ii) constrained, template-based reward shaping significantly improves stability and risk-adjusted returns. Modest gains in episode return translate into large practical improvements in TSC metrics, including up to 28% delay reduction and 9% throughput increase. Our results suggest that LLMs are particularly well-suited as architects of training processes, where their semantic reasoning can structure curricula and rewards for efficient, robust MARL, while leaving execution to domain-appropriate learned controllers.
@@ A.1 Per-Seed Performance Breakdown
Table 4 provides complete per-seed performance data over episodes 180–199. A8 exhibits the widest performance range (11.36 points), with seed 400 achieving the best result (166.08) but seed 500 falling below baseline (154.72). In contrast, A7 shows strong consistency across all four seeds (range: 2.39 points), with all seeds exceeding baseline.
Table 4: Per-seed traffic control performance (episode return over episodes 180–199). Training rewards extracted from simulation logs. All configurations completed 4 seeds each.
@@
A7’s worst-case performance (157.76) exceeds the baseline mean (157.04) by 0.72 points, while A8’s worst-case (154.72) falls 2.32 points below baseline. This represents a critical difference: A7 guarantees improvement regardless of random seed, whereas A8 carries 25% risk of degradation.
@@
  • A2: $\pm$[1.93, 0.29, 2.59, 4.22] points (max deviation: 4.22)
  • A7: $\pm$[0.76, 0.94, 1.28, 1.11] points (max deviation: 1.28)
@@ A.2 Curriculum Adaptation Dynamics
Figure 7: Curriculum difficulty trajectories over 200 episodes for representative seeds. Both A2 and A7 (sharing llm_adaptive mode) exhibit bidirectional adaptation: difficulty increases when performance improves but decreases after performance drops. Shaded regions indicate episodes where difficulty was reduced to prevent overchallenge. Note: A8 uses llm mode with different adaptation characteristics and is not plotted here.
@@ A.4 Computational Cost Breakdown
For a full 4-seed ablation study: $\sim$\$1.88 (A7 with 4 seeds) + \$1.28 (A2 with 4 seeds) + \$1.28 (A8 with 4 seeds) = \$4.44 total.
@@ A.5 Traffic Metrics: Multi-Seed Analysis
Table 7: Traffic efficiency metrics extracted from simulation logs (late-training episodes, seed 200 only). Single-seed results shown; future work will aggregate across all seeds for stronger statistical support.
@@
Current analysis uses seed 200 only due to time constraints. Future work will extract and aggregate traffic metrics across all seeds (seeds 200, 300, 400, 500 for all configurations A2, A7, A8) to provide mean $\pm$ SD for each metric, strengthening statistical rigor.
@@ A.7 Additional Risk Metrics
A2 (Baseline) 152.82 4.22 75% (3/4)
A7 (LLM Rewards) 1.53 2.16 157.76 1.28 100% (4/4)
A8 (Alt Curriculum) 0.67 1.12 154.72 5.51 75% (3/4)
@@ A.8 Configuration Details
  • Reward: LLM-generated with difficulty-adaptive weighting, $w(d)=0.1+0.4\cdot d$ for normalized difficulty $d\in[0,1]$
  • Seeds: 200, 300, 400, 500 (all completed)